OpenStudy (anonymous):

how can I solve this problem? sqrt(2x) +4=x

myininaya (myininaya):

get the sqrt by itself and then square both sides

OpenStudy (anonymous):

yeah, that seems the only way

OpenStudy (anonymous):

and then use the root formula

myininaya (myininaya):

sqrt(2x)=x-4 2x=(x-4)^2 2x=x^2-8x+16 x^2-8x-2x+16=0 x^2-10x+16=0

OpenStudy (anonymous):

darn, beat me to it yeah, jk about the formula

OpenStudy (radar):

(x-8)(x-2)=0

OpenStudy (anonymous):

thank you!

OpenStudy (radar):

x=8 x=2

OpenStudy (radar):

Check out the 2 as it may not work.

myininaya (myininaya):

if you ever have sqrts in your problem, get rid of them by getting just one sqrt on one side of the equation and square both sides to get rid of it. the process will be continued if you have more than one

myininaya (myininaya):

yes always check the x's that you get whenever you raise both sides to an even power which we did to get rid of the sqrt

OpenStudy (radar):

With 2 you are using the negative root.

OpenStudy (radar):

So I guess it is o.k, true you will lose negative sign when squaring so try and keep track.

myininaya (myininaya):

x=2 will not work since sqrt(4) does not equal -2